OpenStudy (jjuden):

Solve each equation, if possible. Write irrational numbers in simplest radical form. Identify the strategy you used to get your solution. 1. x^2 – 9 = 0 2. x^2 + 4x + 2 = 0

OpenStudy (anonymous):

so r they 2 separate ones?

OpenStudy (jjuden):

i dont get this

OpenStudy (jjuden):

yes too separate ones

OpenStudy (anonymous):

1.(x+3)(x-3)

OpenStudy (jjuden):

how did you get that ?

OpenStudy (chris911):

x^2 - 9 = 0 x^2=9 x= 3 or -3

OpenStudy (jjuden):

ok thank you :)

OpenStudy (chris911):

Take the middle term, divide it in half, and square it. (-2)^2=4 That means the third term needs to be 4. Get it there by adding 2 to both sides of the equation. x^2-4x+4=2 Factor (x-2)^2=2 x-2=sqrt2 x=2+sqrt2

OpenStudy (jjuden):

thank you